OpenStudy (anonymous):

find the equation of the tangent to the parabola x^2=16y at the point (4,1)

OpenStudy (callisto):

Differentiate both sides with respect to x, what do you get?

OpenStudy (anonymous):

i got this so far, y=x^2/16 dy/dx=2x/16 = x/8 at (4,1) dy/dx =4/8 =1/2

OpenStudy (callisto):

Yup! So, slope of tangent = m = 1/2 Equation of tangent required is y- y1 = m (x -x1) (x1, y1) = (4,1) Can you do it now?

OpenStudy (anonymous):

yeah but i don't know hw i keep getting x-2y+4

OpenStudy (anonymous):

the answer is x-2y-2=0

OpenStudy (callisto):

y- 1 = (1/2) (x -4) Multiply both sides by 2 2 (y-1) = x-4 Can you continue??

OpenStudy (anonymous):

yeah thanks !!

OpenStudy (callisto):

Welcome :)